Solve the following systems of equations using the substitution method.
What type of system is it? Name the solution if there is one.
. y=-2x
4x-3y=12

Answer by rfer(16322)

4x-3(-2x)=12
4x+6x=12
10x=12
x=1.2
---------------
4*1.2-3y=12
4.8-3y=12
-3y=12-4.8
-3y=7.2
y=7.2/-3
y=-2.4
------------
(1.2,-2.4)
-------------
y=-2x
-2.4=-2*1.2
-2.4=-2.4